# A Unified Framework for N-Dimensional Visualization and Simulation: Implementation and Evaluation including 4D Boolean Operations

### v2

@@ 3.4.4 Dynamic Tessellation via Direct Quickhull
A facet that is found to intersect in the narrow phase is clipped by the (N1)(N-1)-dimensional hyperplane spanned by the opposing facet, and the clipped facet is then tessellated into multiple smaller facets. This entire process is illustrated in Figure 4 using an analogy in 3D. As the figure shows, clipping can decompose the original facet, an (N1)(N-1)-dimensional simplex, into an (N1)(N-1)-dimensional convex polytope with more than NNvertices. Consequently, it is necessary to efficiently re-tessellate the resulting polytope back into multiple smaller facets.
@@
However, in an implementation that assumes floating-point arithmetic, this theoretical classification alone is insufficient. Problems can arise from numerical errors, such as an intersection that should theoretically be a single point being detected as multiple points, or the classification becoming unstable when a vertex is extremely close to the hyperplane. To address this challenge, my implementation introduces a two-stage vertex merging process. First, it unifies points that are in close proximity to each other within the computed set of intersection vertices. Second, it evaluates the distance between each vertex of the merged intersection region and the vertices of the original facet; if they are close, the intersection vertex is completely snapped to the coordinates of the facet vertex. While this operation may introduce minor deformations to the mesh geometry, it plays a crucial role in maintaining the topological consistency of the original facet within the tessellated mesh.
After the merging process, the final tessellation is performed. First, I apply the simplicial tessellation feature of Direct Quickhull to the combined surface-like point cloud of the P-side vertices and the intersection region vertices to tessellate the P-side region into multiple smaller facets. Next, a similar operation is performed for the N-side vertices and the intersection region vertices to tessellate the N-side region. Through these two processes, the original facet is tessellated such that its mesh structure is consistent with the clipping plane. This approach demonstrates that Direct Quickhull can effectively use its capabilities even in dynamic mesh processing, thereby guaranteeing the $N$-dimensional extensibility of the entire algorithm.
@@ 3.4.5 Inside-Outside Testing and Result Construction
To construct the final result of a Boolean operation—namely, the union, intersection, or difference from the set of tessellated facets, it is necessary to classify whether each facet is inside or outside the opposing object. For this inside-outside test, my implementation adopts the ray casting method. The fundamental principle of this method is based on intersection parity. A ray is cast from the centroid of the facet under test toward the opposing surface, and the number of intersections is counted. The state is then determined according to whether this number is even or odd, and the same logic applies in $N$-dimensional space. In my implementation, the ray is parallel to the x-axis, and its endpoint is set to a point slightly beyond the opposing mesh’s AABB. This ensures both computational efficiency and stability. To make this method operate robustly in a floating-point environment, it is necessary to handle degenerate cases where the ray passes through a facet’s vertices or edges, as shown in Figure 6. In an implementation that introduces a tolerance, an intersection with a vertex or edge, which should theoretically be a single point of intersection, can be detected as multiple, slightly different intersection points with the various facets that share that feature. This disrupts the parity of the intersection count and causes the inside-outside test to fail. To prevent this misclassification, my implementation employs a two-stage solution. First, the coordinates of the detected intersection points are discretized based on a tolerance, and multiple proximate intersection points are grouped into a single ’intersection event’ so they can be handled uniformly in subsequent processing. Next, the nature of this grouped intersection event is determined. Specifically, it is necessary to distinguish whether the ray is piercing the surface or merely touching it.
@@
To do this, I evaluate the dot product of the ray’s direction vector and the normal vector of every facet included in the event. If the signs of the dot products are biased to one side (either all positive or all negative), the event is considered to be piercing the surface, and the intersection count is incremented by one. If, on the other hand, the signs are distributed between both positive and negative, the ray is determined to be just tangent to the surface, and it is not included in the intersection count. Based on this inside-outside test executed through this series of processes, the tessellated facets are classified according to the type of Boolean operation, and the final mesh is constructed.
@@
However, in cases where the ray is coplanar with a facet, such as the ”Co-planar Facet” case shown in Figure 5, the dot product of the ray and the normal vector is theoretically zero, causing the sign-based test to fail robustly. Therefore, there is no guarantee that intersections can always be correctly determined under these conditions. A practical approach would be to avoid such degenerate geometric configurations by using existing techniques, such as ray perturbation.
3.5 Visualization of N-Dimensional Objects
3.5.1 Cross-Sectioning as the Basic Strategy
This component is responsible for generating and rendering 3D cross-sections based on $N$-dimensional mesh data. The basic strategy to achieve this transformation is based on the concept of sequentially reducing the dimension. This is a process of recursively clipping an $N$-dimensional object with an (N1)(N-1)-dimensional hyperplane, and then clipping the resulting cross-section with an (N2)(N-2)-dimensional hyperplane. In this paper, I refer to this method of sequentially reducing dimensions as Hierarchical Cross-Sectioning. With this approach, the complex problem of computing cross-sections in arbitrary dimensions can always be reduced to the simpler problem of the intersection between an $N$-dimensional facet and an (N1)(N-1)-dimensional hyperplane. In the following sections, as a concrete application of this Hierarchical Cross-Sectioning for $N=4$, I will detail the specific algorithm for generating 3D cross-sections from 4D objects and discuss its implementation challenges.
3.5.2 Facet Clipping Algorithm in the 4D Implementation
To generate a 3D cross-section of an object in 4D space, my implementation computes the geometric intersection between each facet of the 4D object—a tetrahedron—and an axis-parallel hyperplane $w=w_{\text{slice}}$ (where $w=w_{\text{slice}}$ is an arbitrary constant). This procedure is executed independently for every facet of the input mesh according to the following steps.
First, I determine on which side of the clipping hyperplane each of the facet’s four vertices is located. If all vertices are on the same side, the facet does not intersect the hyperplane and skip intersection computation. If the vertices straddle the hyperplane, the intersection point of each constituent edge of the facet with the hyperplane is calculated via linear interpolation. Let the two vertices defining an edge be
The interpolation parameter ttfor determining the position of the intersection point is then calculated from the $w$-coordinates of both vertices and the $w$-coordinate of the clipping plane, $w_{\text{slice}}$, as shown in Equation (6).
$t=\frac{w_{\text{slice}}-w_{0}}{w_{1}-w_{0}}$ (6)
$\mathbf{p}_{\text{isect}}=\mathbf{p}_{1}+t(\mathbf{p}_{1}-\mathbf{p}_{0})$ (7)
The 3D coordinates, obtained by extracting the $x,y,z$ components of the resulting 4D coordinate vector $\mathbf{p}_{\text{isect}}$, directly become the coordinates of a vertex that constitutes the cross-section.
By performing this operation for all edges that intersect the hyperplane, I obtain the set of vertices of the cross-section, which are then connected to form a polygon. The shape of the cross-section generated by the intersection of a single 4D facet and a 3D hyperplane is, within 3D space, either a triangle with three vertices or a convex polygon with four vertices. A 4-vertex convex polygon can be divided into two triangles to obtain a set of triangular polygons for rendering.
3.5.3 Simplicial Decomposition of the Cross-Section
Figure 7: Dimensional dependence of the number of cross-section vertices generated during hyperplane clipping. When clipping a facet in 3D, the number of vertices is always two. However, when clipping a 3D facet in 4D, the number varies between three and four depending on the relative position of the hyperplane, and is thus not uniquely determined. Regarding the color coding in this figure, the blue plane represents the clipping hyperplane, and the red area indicates the portion of it that exists on the dimension being clipped. The darker shades are highlights for visual clarity.
The cross-section generated by the above clipping algorithm is generally an (N2)(N-2)-dimensional convex polytope. For subsequent processing, this must be decomposed into a set of (N2)(N-2)-dimensional simplices. In the 4D implementation, the cross-section is a 2D convex polygon (either a triangle or a quadrilateral) with a maximum of four vertices, as shown in Figure 7. Although the order of the vertex list is undefined, the polygon can be safely tessellate into two triangles by sorting the vertices by angle around the polygon’s centroid and then systematically partitioning it. my implementation efficiently performs this sort using only vector operations, avoiding computationally expensive trigonometric functions.
However, this angle-sorting method cannot be generalized to five or more dimensions $(N\geq 5)$. The maximum number of vertices of an (N2)(N-2)-dimensional cross-section, obtained by clipping an $N$-dimensional facet (which has NNvertices) with a hyperplane, is given by the well-known result in combinatorics shown in Equation (8).
For example, in 5D, the cross-section can be a 3D convex polytope with a maximum of $\lfloor 5/2\rfloor\cdot\lceil 5/2\rceil=6$ vertices. The simplicial tessellation of such a complex polytope is a non-trivial computational problem. As a dimension-independent and general solution to this high-dimensional extensibility challenge, I can use the simplicial tessellation feature of Direct Quickhull, which was detailed in Section 3.3.3. By providing the set of vertices that constitute the cross-section as input to Direct Quickhull, it is possible to efficiently tessellate the interior of the cross-sectional polytope itself into a set of (N1)(N-1)-dimensional facets. As mentioned in Section 3.3.1, my implementation, which omits the interior point search, is well-suited for processing this type of surface-like point cloud. This ensures that the visualization pipeline of my framework is, in principle, extensible to higher dimensions.
3.5.4 Robust Method for Determining Cross-Section Normals
When generating a 3D cross-section for rendering from a 4D object, it is necessary to determine the winding order of the vertices that form the 3D facet. The clipping algorithm from the previous section produces a list of vertices in an undefined order. If this unordered vertex list is used as is, a serious problem arises: many rendering pipelines determine the front and back of a face based on the input order, which would result in inconsistent face orientations and prevent the correct rendering of the shape.
To solve this problem, my method utilizes the normal vector of the original 4D facet as a reference for the correct orientation. Specifically, I first project the original 4D normal vector into 3D space by extracting its $x,y,z$ components, and define this as the true normal vector that the 3D facets should follow. Next, I compute a temporary normal vector from the generated vertex set using the cross product and evaluate its dot product with the true normal vector. If the sign of this dot product is negative, the winding order of the vertices is reversed. This ensures that all triangles constituting the cross-sectional mesh have a consistent vertex winding order that can be correctly interpreted by a general-purpose renderer.
3.6 Pose Representation and Control Scheme for Interactive Exploration
@@ 3.6.1 Pose Representation and Management with Geometric Algebra Rotors
In representing the pose of an $N$-dimensional object, conventional 3D methods like Euler angles or quaternions cannot be directly applied due to issues such as gimbal lock and difficulties in extending them to higher dimensions. To solve this challenge, my framework adheres to the mathematical foundation of Geometric Algebra, as mentioned in Section 3.2, and represents the object’s pose uniformly using a Rotor.
@@
In my implementation, instead of having the user manipulate the rotor directly, I allow them to specify the amount of rotation for each of the six fundamental rotation planes in 4D space $(xy,xz,xw,yz,yw,zw)$ as individual parameters, similar to Euler angles. These parameter values are then composed in a predefined order to internally construct the final rotation operator, the rotor. This design, which separates the intuitive parameters provided to the user from the mathematically robust operator used internally, completely avoids gimbal lock and is easily extensible to higher dimensions. Furthermore, my implementation fixes the clipping hyperplane to be axis-parallel $(w=\text{const})$ and transforms the object’s pose instead. From the user’s perspective, this approach is visually equivalent to clipping a stationary object with an arbitrarily oriented hyperplane. However, it replaces the computationally expensive problem of arbitrary hyperplane clipping with a combination of an efficient pose transformation and an axis-parallel clip, enabling interactive observation of the cross-sections.
3.6.2 Camera-Based Coordinate Transformation
@@
To intuitively understand the geometric essence of this transformation, I consider the 2D space shown in Figure 8 as an example. I denote the position of a target vertex $\mathbf{P}$ in the world coordinate system as $\mathbf{P}^{(\mathrm{W})}$ and the position of the camera $\mathbf{C}$ as $\mathbf{C}^{(\mathrm{W})}$. The coordinate transformation takes $\mathbf{P}^{(\mathrm{W})}$, $\mathbf{C}^{(\mathrm{W})}$, and the rotor $R$ representing the pose as input. The transformation process first derives the basis vectors that span the camera coordinate system, $\mathbf{e}_{\text{right}}^{(\mathrm{W})}$ and $\mathbf{e}_{\text{fwd}}^{(\mathrm{W})}$, from the rotor $R$ representing the camera’s pose, as shown in Equation (11). Next, the relative position vector $\mathbf{v}^{(\mathrm{W})}$ from the camera’s position to the object’s vertex is calculated as shown in Equation (12).
@@
This entire sequence of processes can be interpreted as two separate geometric operations. Equation (12) corresponds to a translation that aligns the origin of the world coordinate system with the camera’s position. Equation (13) corresponds to a rotation that aligns the coordinate system with the camera’s orientation. By using homogeneous coordinates, these two operations can be combined into a single $(N+1)\times(N+1)$ matrix, analogous to the view matrix commonly used in computer graphics. The description using dot products in this paper is specifically intended to directly illustrate the geometric meaning of the rotation operation.
3.6.3 Design of the Control Scheme for Interactive Exploration
To address the challenge of mapping the six independent rotational degrees of freedom required to control the camera’s orientation in 4D space onto a standard mouse and keyboard, I adopted two design principles: orthogonality of controls and leveraging existing control conventions (specifically, those of First-Person Shooters, or FPS). The 3D translations (WASD keys) and pitch rotation (mouse vertical movement) follow standard conventions. The primary challenge, which is the assignment of a rotation to the horizontal mouse movement ($x$-axis), is resolved by a modal input switch using a modifier key (Left Control). In its normal mode, horizontal mouse movement controls the primary yaw rotation ($xz$-plane) for 3D exploration. While the modifier key is held down, this input is switched to control the rotation in the $xw$-plane, which serves a similar role in 4D exploration. This design ensures that a single input always corresponds orthogonally to a single operation, resulting in a highly predictable control scheme that allows the user to operate while being clearly conscious of switching between 3D and 4D exploration modes.
3.6.4 Extension of the Manipulation System to N Dimensions
This design principle can be generalized to $N$-dimensions. I apply a systematic approach to address the rotational degrees of freedom in an $N$-dimensional space, which increase quadratically as N(N1)/2N(N-1)/2. First, translations and rotations within the 3D cross-section that the user directly observes are kept on the standard control scheme (WASD/mouse), regardless of the total number of dimensions. This provides the user with a consistently stable operational base. Next, for the remaining (N(N1)/2)3(N(N-1)/2)-3 rotations outside the observed space, I consider an approach that generalizes the design of my 4D implementation. Specifically, horizontal mouse movement is consistently assigned to rotations that occur outside the observed space; the user selects the target high-dimensional axis $(w,u,\ldots)$ via a combination of modifier keys, which in turn manipulates the corresponding rotation plane $(xw,xu,\ldots)$. Because this approach dedicates the primary rotational input to high-dimensional exploration while allowing the control scheme to be systematically extended, it is expected to provide a scalable foundation for exploring $N$-dimensional space while preserving control orthogonality.
3.7 Demonstration of Extensibility
The core of this framework’s extensibility lies in the design principle of separating topology and geometry. This section demonstrates the effectiveness of this design through a case study in which a 4D non-rigid body physics simulation is integrated.
A practical challenge in this integration is the mismatch of data formats. For computational efficiency, the physics simulation requires unique, non-duplicated vertex data. The rendering pipeline, on the other hand, requires duplicated vertex data to accommodate correct normals and UVs. The separated architecture of my framework resolves this discrepancy through index mapping. During initialization, the physics engine extracts the set of unique vertices from the rendering mesh’s geometry and builds a map, just once, that maintains the correspondence between the indices of the unique vertices and the indices of the original, duplicated vertices. During the simulation, computations are performed only on the set of unique vertices. Then, in each frame, this map is used to efficiently update the entire vertex array for rendering.
This pattern—maintaining an internally optimized data format and converting it to the required format only during communication is a general solution for integrating any external component that requires a different data format, and is not limited to this case study. This demonstrates that the design principle of separating topology and geometry is what enables the high modularity and extensibility of my framework.
4 Plex: An N-Dimensional Mesh Data Exchange Ecosystem

### v3

@@ 3.4.3 Dynamic Tessellation via Direct Quickhull
A facet that is found to intersect in the narrow phase is clipped by the (N1)(N-1)-dimensional hyperplane spanned by the opposing facet, and the clipped facet is then tessellated into multiple smaller facets. As illustrated in Fig. 4, a facet that is originally an (N1)(N-1)-dimensional simplex may, through cutting, be decomposed into an (N1)(N-1)-dimensional convex polytope whose number of vertices exceeds $N$. In such cases, the polytope after cutting must be subdivided into multiple smaller facets.
@@
After classifying the vertices of a facet, the Direct Quickhull simplex-filling procedure is applied to a point set that combines the intersection points between the facet and the cutting hyperplane with the vertices on the positive (P) side, thereby re-subdividing the P-side region into multiple smaller facets. The same operation is then applied to the intersection points together with the vertices on the negative (N) side, and the N-side region is re-subdivided. Through these two processes, the facet is re-subdivided so as to be consistent with the cutting plane.
In the 4D implementation, because floating-point arithmetic is assumed, numerical errors may cause issues such as detecting multiple intersection points where a single point should theoretically exist, or unstable classification when facet vertices lie extremely close to the cutting hyperplane. To address this issue, the following merge procedure was introduced.
First, the intersection point set between the cutting hyperplane and the facet is computed, and duplicates are removed based on a quantization key. Next, among the facet vertices, points that are extremely close to the cutting hyperplane are searched. Let the facet vertex be $\mathbf{v}$, the normal of the cutting hyperplane be $\mathbf{n}$, an arbitrary point on the cutting hyperplane be $\mathbf{p_{\text{0}}}$, and the tolerance be $\varepsilon$. A point that is extremely close to the cutting hyperplane is defined as a point satisfying Eq. (5).
$\lvert(\mathbf{p_{\text{0}}}-\mathbf{v})\cdot\mathbf{n}\rvert\leq\varepsilon$ (5)
If there exists any vertex satisfying Eq. (5), the set consisting of the intersection point set and those vertices is treated as the intersection candidate set, and duplicates are removed again based on the quantization key. When an intersection point and a facet vertex are classified under the same quantization key, the intersection coordinates are replaced with the facet vertex coordinates. Although this operation may introduce minor deformation of the mesh, it is positioned as a process to suppress gaps and overlaps in the subdivided mesh.
@@ 3.4.4 Inside-Outside Testing and Result Construction
To construct the final result of a Boolean operation (union, intersection, or difference) from a set of tessellated facets, it is necessary to classify whether each facet lies inside or outside the other object. For inside-outside classification, this architecture adopts a ray-casting method, which is used in Boolean operations that emphasize interactivity [4]. This method determines whether a point lies inside or outside an object based on the parity of the number of intersections between a ray cast from the point and the surface of the target object.
In the 4D implementation, for the sake of implementation simplicity, the centroid of the facet under consideration is used as the ray origin, and the ray is cast parallel to the $x$-axis. The ray endpoint is set to a point slightly beyond the axis-aligned bounding box (AABB) of the opposing mesh. As illustrated in Fig. 6, in degenerate cases where the ray passes through a facet vertex or edge, floating-point arithmetic may cause intersections that should theoretically be a single event at a vertex or edge to be detected as multiple, slightly different intersection points associated with multiple facets sharing that vertex or edge. This can lead to misclassification in inside-outside tests. To address this issue, we quantize the detected intersection points using a tolerance and group points with the same quantization key into a single intersection event. Next, the nature of this grouped intersection event is determined. Specifically, it is necessary to distinguish whether the ray is piercing the surface or merely touching it.
@@
To do this, we evaluate the dot product of the ray’s direction vector and the normal vector of every facet included in the event. If the signs of the dot products are biased to one side (either all positive or all negative), the event is considered to be piercing the surface, and the intersection count is incremented by one. If, on the other hand, the signs are distributed between both positive and negative, the ray is determined to be just tangent to the surface, and it is not included in the intersection count. Based on this inside-outside test executed through this series of processes, the tessellated facets are classified according to the type of Boolean operation, and the final mesh is constructed.
@@
However, in cases where the ray lies on the same plane as a facet, such as the “co-planar facet” case illustrated in Fig. 5, the dot product between the ray direction and the facet normal vector is theoretically zero. As a result, misclassification may occur in sign-based decision methods. Therefore, under such conditions, there is no guarantee that intersections can always be classified correctly. To avoid these degenerate geometric configurations, techniques such as ray perturbation can be considered as candidate approaches.
3.5 Coordinate Transformations for High-Dimensional Exploration
@@ 3.5.1 Pose Representation and Management with Geometric Algebra Rotors
In this architecture, the orientation of an $N$-dimensional object is represented using rotors, in accordance with geometric algebra, which serves as the mathematical foundation described in Section 3.1.
@@
In the 4D implementation, users do not manipulate rotors directly; instead, rotation amounts are specified as individual parameters, analogous to Euler angles, for each of the six fundamental rotation planes in 4D space $(xy,xz,xw,yz,yw,zw)$. These parameter values are composed in a prescribed order to internally construct a rotor as the final rotation operator, which is then used for rotation computation.
3.5.2 View Transformation
@@
To intuitively understand the geometric essence of this transformation, we consider the 2D space shown in Figure 7 as an example. We denote the position of a target vertex $\mathbf{P}$ in the world coordinate system as $\mathbf{P}^{(\mathrm{W})}$ and the position of the camera $\mathbf{C}$ as $\mathbf{C}^{(\mathrm{W})}$. The coordinate transformation takes $\mathbf{P}^{(\mathrm{W})}$, $\mathbf{C}^{(\mathrm{W})}$, and the rotor $R$ representing the pose as input. In the transformation process, the rotor $R$ representing the camera orientation and the basis vectors of the world coordinate system are first used to rotate the basis vectors according to Eq. (8), thereby deriving the basis vectors $\mathbf{e}_{\text{right}}^{(\mathrm{W})}$ and $\mathbf{e}_{\text{fwd}}^{(\mathrm{W})}$ that span the camera coordinate system. Next, the relative position vector $\mathbf{v}^{(\mathrm{W})}$ from the camera’s position to the object’s vertex is calculated as shown in Equation (9).
@@
This entire sequence of processes can be interpreted as two separate geometric operations. Equation (9) corresponds to a translation that aligns the origin of the world coordinate system with the camera’s position. Equation (10) corresponds to a rotation that aligns the coordinate system with the camera’s orientation. By using homogeneous coordinates, these two operations can be combined into a single $(N+1)\times(N+1)$ matrix, analogous to the view matrix commonly used in computer graphics. The dot-product-based description in this paper is intended to make the geometric meaning of the rotation operation explicit.
3.6 Visualization of N-Dimensional Meshes
3.6.1 Cross-Sectioning as the Basic Strategy
This component generates and renders 3D cross-sections from $N$-dimensional mesh data. The basic strategy for realizing this process is based on the idea of sequentially reducing dimensionality. This is a process of recursively clipping an $N$-dimensional object with an (N1)(N-1)-dimensional hyperplane, and then clipping the resulting cross-section with an (N2)(N-2)-dimensional hyperplane. In this paper, we refer to this method of sequentially reducing dimensions as Hierarchical Cross-Sectioning. With this approach, complex cross-section computation problems in arbitrary dimensions are always reduced to a simpler problem: the intersection between an $N$-dimensional facet and an (N1)(N-1)-dimensional hyperplane.
However, because the information obtained through sequential hyperplane slicing is limited to cross-sections, it involves information loss, and the original shape cannot be uniquely reconstructed from the cross-sections alone. A quantitative evaluation of such perceptual limitations is beyond the scope of this paper.
In the following sections, as a concrete application of this Hierarchical Cross-Sectioning for $N=4$, we will detail the specific algorithm for generating 3D cross-sections from 4D objects and discuss its implementation challenges.
3.6.2 Facet Clipping Algorithm in the 4D Implementation
To generate 3D cross-sections of objects embedded in 4D space, the present implementation computes the geometric intersections between each facet of the 4D mesh, represented as tetrahedra, and an axis-aligned hyperplane $w=w_{\text{slice}}$, where $w_{\text{slice}}$ is an arbitrary constant. This computation is performed for all facets of the input mesh. Cross-section generation based on this type of slicing plane has also been adopted in existing studies on interactive 4D visualization [5]. In this implementation, no mechanism is provided to arbitrarily specify the orientation of the cutting hyperplane; cross-sections are always generated using the axis-aligned plane $w=w_{\text{slice}}$. When observing the cross-sections from arbitrary directions, changes in the viewing direction are handled as view transformations.
First, we determine on which side of the clipping hyperplane each of the facet’s four vertices is located. If all vertices are on the same side, the facet does not intersect the hyperplane, and we skip the intersection computation. If the vertices straddle the hyperplane, the intersection point of each constituent edge of the facet with the hyperplane is calculated via linear interpolation. Let the two vertices defining an edge be
The interpolation parameter $t$ for determining the position of the intersection point is then calculated from the $w$-coordinates of both vertices and the $w$-coordinate of the clipping plane, $w_{\text{slice}}$, as shown in Equation (12).
$t=\frac{w_{\text{slice}}-w_{0}}{w_{1}-w_{0}}$ (12)
$\mathbf{p}_{\text{isect}}=\mathbf{p}_{0}+t(\mathbf{p}_{1}-\mathbf{p}_{0})$ (13)
The 3D coordinates obtained by extracting the $x,y,z$ components of the resulting 4D coordinate $\mathbf{p}_{\text{isect}}$ are directly used as the vertex coordinates constituting the cross-section.
By performing this operation for all edges of facets that intersect the hyperplane, the vertex set of the cross-section is obtained, and polygons are constructed by connecting these vertices. The shape of the cross-section generated by the intersection of a single 4D facet and a 3D hyperplane is, within 3D space, either a triangle with three vertices or a convex polygon with four vertices. For a convex polygon with four vertices, it is divided into two triangles to obtain a set of triangular polygons for rendering.
3.6.3 Simplicial Decomposition of the Cross-Section
Figure 8: The dimensional dependence of the number of cross-section vertices generated by hyperplane cutting is as follows. In three dimensions, facet cutting produces two vertices, whereas in four dimensions the number of vertices becomes three or four depending on the hyperplane’s relative position. Regarding the color coding in this figure, the blue plane represents the clipping hyperplane, and the red area indicates the portion of it that exists on the dimension being clipped. The darker shades are highlights for visual clarity.
The cross-sections generated by the aforementioned cutting algorithm form an (N2)(N-2)-dimensional convex polytope in $N$ dimensions. For subsequent processing, this must be decomposed into a set of (N2)(N-2)-dimensional simplices. The maximum number of vertices of an (N2)(N-2)-dimensional cross-section obtained by cutting an $N$-dimensional facet (with $N$ vertices) by a hyperplane is given by  (14).
For example, in 5D, the cross-section of a facet can be a 3D convex polytope with a maximum of $\lfloor 5/2\rfloor\cdot\lceil 5/2\rceil=6$ vertices. For simplicial decomposition of such polytopes, the simplex-filling function of Direct Quickhull described in Section 3.3.3 is used. By providing the vertex set constituting the cross-section as input to Direct Quickhull, the cross-section is subdivided into a set of (N2)(N-2)-dimensional simplices.
In the 4D implementation, as illustrated in Fig. 8, the cross-section is a two-dimensional convex polygon with at most four vertices. Therefore, the vertex order is arranged by angle sorting using the cross product, and a four-vertex two-dimensional convex polygon is subdivided into two triangles.
3.6.4 Determination of Cross-Section Normals and Vertex Winding Order
In the 4D implementation, when generating 3D cross-sections to be rendered from a 4D object, it is necessary to determine the vertex winding order of the 3D facets. The cutting algorithm described in the previous section produces an unordered set of vertices. If this unordered vertex set is used as is, many rendering pipelines determine front-facing and back-facing polygons based on the input order, which can result in inconsistent face orientations and, consequently, incorrect rendering of the shape.
To address this issue, the present implementation uses the normal vector of the original 4D facet as a reference for the correct orientation. First, the $x,y,z$ components of the 4D facet normal vector are extracted to project the normal vector into 3D space, and this projected vector is defined as the reference normal vector for determining the winding order of the 3D facet. Next, the dot product between this reference normal vector and a provisional normal vector computed by the cross product from the vertex set of the generated 3D facet is evaluated. If the sign of the dot product is negative, the vertex winding order is reversed, thereby ensuring consistent winding of the triangles constituting the cross-section mesh.
3.7 Interaction Scheme for Interactive Exploration
3.7.1 Interaction Scheme Design for the 4D Implementation
In the 4D implementation, among the six rotational degrees of freedom, the operations handled by the mouse and keyboard are assigned according to the following policy. First, to keep the mapping between inputs and actions as orthogonal as possible, we adopt FPS navigation as the basis, separating translation to the keyboard (WASD) and viewpoint rotation to the mouse. Navigation centered on continuous first-person movement has been evaluated in comparative studies of interaction techniques for exploration tasks [21, 16].
Second, to handle high-degree-of-freedom rotations with limited 2D input, the assignment of horizontal mouse movement ($x$-axis) is switched using a modifier key (LeftControl). In the default state, it is assigned to yaw rotation (the $xz$ plane), and while the modifier key is held down, it is switched to rotation in the $xw$ plane. Such a “spring-loaded mode” (quasimode), which is active only while the key is held, has been discussed as a design that can reduce mode errors compared to persistent modes [11].
3.7.2 Application of Interaction Design Policies to N Dimensions
In this section, we describe an interaction scheme obtained by applying the assignment policy used in the aforementioned 4D implementation to $N$ dimensions. Since the rotational degrees of freedom in an $N$-dimensional space increase to N(N1)/2N(N-1)/2, translation and rotation within the observed 3D cross-section are designed to use dimension-independent operations (WASD/mouse). In contrast, for rotations outside the observed space, following the same principle as in the 4D implementation, additional degrees of freedom are introduced incrementally while preserving input orthogonality. Specifically, horizontal mouse movement is assigned to rotations outside the observed space, higher-dimensional axes $(w,u,\ldots)$ are selected via combinations of modifier keys, and the corresponding rotation planes $(xw,xu,\ldots)$ are designated as the manipulation targets. This type of switching is consistent with a design policy in which constraints are used to control the degrees of freedom handled by the user [20].
3.8 Design Based on Topological and Geometric Separation
In physical simulations and related applications, when handling geometry that changes over time while sharing the same topology, a unique vertex array constituting the mesh is typically assumed as input. In contrast, visualization and rendering pipelines treat facets as the minimal unit; therefore, when a collection of facets is represented using vertex sets, duplicate indices arise, and such pipelines assume vertex arrays that include these duplicates as input. As a result, the data formats required by update systems, which demand unique vertex arrays, and by visualization and rendering systems, which assume vertex arrays containing duplicated indices, do not coincide. Under the assumptions of the update and visualization/rendering systems considered in this study, it is therefore not possible to integrate both within a single architecture without separating topology and geometry.
To address this issue, the proposed architecture adopts a design in which, at the initialization stage, a correspondence table (index mapping) between a unique vertex array and a vertex array containing duplicated indices is maintained, derived from the mesh data.
This design allows update systems that take a unique vertex array as input and visualization/rendering systems that assume vertex arrays with duplicated indices to be handled simultaneously within the same architecture. Update processing is completed solely on the unique vertex array, while reflection to visualization and rendering is separated as a transfer process based on the correspondence table.
In the 4D implementation, based on this design, a non-rigid body simulation based on XPBD is integrated, and it is confirmed that mesh vertex coordinates are updated over time and reflected in cross-section generation and rendering. However, evaluation of the validity or accuracy of the physical model is not included in the scope of this study.
4 Plex: An N-Dimensional Mesh Data Exchange Format